OpenStudy (anonymous):

f(x) = 3x^2 - 6 __________ x^ + 2x h(x) = 6x + 2 Find: f(h(-0.5)) ( ---- means its an fraction) Please helpp !!!

OpenStudy (anonymous):

we first, find \[h(-0.5)=?\]

OpenStudy (anonymous):

so, what do you get?

OpenStudy (anonymous):

i dont get it at all i tried i dont get the answer thats one of the choices

OpenStudy (espex):

@amandaschlor you solve h(x) using -0.5 as your input, then you take the output from that and plug it into your f(x) as its input, once you have solved f(x) you will have your answer.

OpenStudy (anonymous):

input & output?

OpenStudy (espex):

A function is like a machine, you give it "input" and it spits out "output". So whatever you plug in for 'x' is your input, and the results from the math is your output.

OpenStudy (espex):

They ask you to solve f(h(-0.5)), so you start by feeding -0.5 into h(x). h(x) = 6x + 2 => h(-0.5)=6(-0.5)+2

OpenStudy (anonymous):

so then what would you do?

OpenStudy (espex):

Solve that and plug the answer in for each 'x' in your f(x) equation.

OpenStudy (anonymous):

just find h(-0.5) first. dont check the answer yet

OpenStudy (anonymous):

ploug in -0.5 for x in the equation for h(x) what do you get?

OpenStudy (anonymous):

i dont get any of the choice answer, the choices are, 10,44,3,7 && i got 5?

OpenStudy (anonymous):

like I said, do not check with the answer yet. just do it step by step what value do you get for h(-0.5)?

OpenStudy (anonymous):

i am just so confused i dont get any of this.

OpenStudy (anonymous):

good. so, do this step find h(-0.5) = ?

OpenStudy (anonymous):

plug in x=-0.5 in the equation for h(x)

OpenStudy (anonymous):

okay i did it equals -1

OpenStudy (espex):

Now take your -1 and plug it for every 'x' in your equation for f(x)\[\frac{3x^{2} - 6}{x^{2} + 2x}\] to get your final answer.

OpenStudy (anonymous):

\[ f(h(-0.5))=f(-1)=\frac{3(-1)^2-6}{(-1)^2+2(-1)}=? \]